OpenStudy (anonymous):

solve 5x+y=3 2x-3y=25 by elimination techniche

OpenStudy (anonymous):

5x+y=3 multiply this equation by 3 so we can eliminate the -3y below this gives us: 15x+3y=9 2x-3y=25 ----------- = 17x+0y=34 ==> x=34/17=2 Insert x=2 into any of the two original equations and you get: equation 1: 5*2+y=3 <==> y=-7 equation 2: 2*2 -3y=25 <==> -3y=21 <==> y=-7 So you get the answer x=2 and y=-7 You can check that the values of x and y are correct by inserting them to any of the two equations. Hope this helps!?! //Ali